OpenStudy (anonymous):

there are 20 coins in total. 48 cents is the total amount, how many coins are pennies and how many coins are nickels.

OpenStudy (texaschic101):

n + p = 20 --> n = p - 20 .05n + .01p = .48 now sub p - 20 in for n and solve for n once n is found sub that into 1st equation and solve for p you can then check your answer by subbing in known values into the 2nd equation and if they come out equal you are correct, however, if they do not come out equal, then there has been a mistake made.